Status Report

NASA Space Station On-Orbit Status 16 July 2004

By SpaceRef Editor
July 17, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 16 July 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally except those noted previously or below.  

Update on SM cooling pump failure:   The failed pump panel of the Russian segment’s internal cooling system was successfully replaced with a new unit.  After reviewing a five-page list of instructions uplinked overnight, CDR Gennady Padalka deinstalled the failed SPN pump panel at the 4SPN1 work site of the Service Module (SM)’s no. 2 internal cooling loop (KOB-2) and replaced it with a spare of new design, delivered on Progress 13P.  The new pump panel was then tested successfully from the controlling laptop, and KOB-2 returned to nominal operation.    [Each of the two thermal loops (KOB-1 & KOB-2) has two pump panels, each of which has two pumps hard welded to it.  The automatic command algorithm does not allow a panel to run on only one of them, thus failure of one pump effectively fails the entire panel.  The replacement today used a new design with individually replaceable pumps, so that a pump failure will not require swap-out of the entire SPN.  The other SPNs will in the future also be replaced with the new design.]

Update on US EMU spacesuit troubleshooting:  A 27-page procedure for removing, disassembling and inspecting the water pump of EMU (extravehicular mobility unit) 3013 was uplinked overnight.  The task, estimated to take 4 hrs of crewtime, is scheduled for Mike Fincke on Monday (7/19).  EMU 3005 meanwhile will retain its integrity for the time being.  Today, Mike was provided with a CBT (computer-based training) video for review over the weekend.  He also had time scheduled for printing out the procedural instructions and gathering the tools required for the EMU activity.   [Monday’s round of troubleshooting is intended to determine what might be keeping the impeller of the suit’s water pump from rotating.  One possible cause could be particle contamination jammed in the very tight radial clearance of the rotor, but there may be other things as well, which Mike’s investigation hopefully will reveal.  The impeller is connected to the electric pump motor by a magnetic coupling.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Update on Russian software upgrading:   In support of the ongoing troubleshooting of the failed uploading of the new Russian SM 7.02 software (reported yesterday), CDR Padalka today downloaded data dumps (images) from the SM’s TVM (Terminal Computer) and TsVM (Central Computer) system to the Central Post computer #1 (KTsP-1) for review on the ground.  As currently planned, Moscow will re-attempt the 7.02 upload to TVM and TsVM next Tuesday (7/20) morning.

The crew spent several hours to continue the time-consuming task of stowing excessed equipment and trash in Progress-249/14P for disposal.  As of last night, they had completed prepacking 37 of the 39 items on the 14P US trash packing list.   [Food containers will be gathered and stowed closer to 14P hatch closure time.  The only item not expected to be ready for disposal is the failed Node LHA (lamp housing assembly), which still needs to be removed.  If there isn’t enough time for that prior to 14P undock, it will become a trash candidate for 15P.] 

Both crewmembers in turn took their third periodic On-Orbit Hearing Assessment (O-OHA) test, a NASA environmental health systems examination to assess the efficacy of acoustic countermeasures.   [The O-OHA audiogram test involves minimum audibility measurements for each ear over a wide range of frequencies and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter).  To conduct the testing, the experimenter is supported by special “EarQ” software on the MEC (medical equipment computer).  The baseline test is required not later than about Flight Day 14 for each new Expedition and is then performed once per month.]

FE/SO Fincke performed the regular once-a-week maintenance reboot on the two operational PCS laptops and the restart of the OCA comm router laptop (every two weeks).

Gennady conducted the routine maintenance of the SM’s life support system (SOZh) and also prepared the daily IMS “delta” file update, which had been added to the discretionary Russian task list.

At the Lab RWS (robotics work station), Mike disconnected the UOP-DCP (utility outlet panel-to-display & control panel) bypass power cable used yesterday for the MSS (mobile service system) video camera characterization.   [During the ground-commanded activity, it was discovered that one of the SSRMS (Space Station Remote Manipulator System)’s tip elbow lights is not working (the second one is OK).  This could be due to either a circuitry/logic problem or a burnt-out bulb.  Full redundancy can be restored by swapping out the camera with an existing on-board spare during a future EVA.]

At 2:30pm EDT, Mike Fincke is scheduled to discuss onboard experiment operations in a 15-min. payload teleconference with the ISS Program Scientist (Don Thomas), Increment 9 Lead Scientist (Janice Voss) and Increment 9 Lead Payload Operations Director (POD, Lamar Stacy).

The crew completed their daily 2.5-hr. physical workout program on TVIS treadmill, RED resistive exerciser and VELO ergometer with load trainer, and Mike performed the periodic (every other week) inspection of the RED.

Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Today’s CEO (crew earth observation) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Patagonian Glaciers (unusual clearing on the Pacific side of the southern Andes.  Crew was asked to shoot any small glacier tongues near nadir), Sao Paulo, Brazil (nadir pass over this gargantuan city.  Margins stretch left and right of track), and Caracas, Venezuela (nadir pass).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

U.S. & Russian Segment Status  (as of today, 1:17pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (16A, =lowest setting).  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 752; temperature (deg C) — 24.6; ppO2 (mmHg) — 164.4; ppCO2 (mmHg) — 3.4.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 19.3.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 21.3.
  • Node:  Pressure (mmHg) — 747.7; temperature (deg C) — 22.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 749.9; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 750.1; temperature (deg C) — 24.2; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Dual angle/”blind” mode, non solar-tracking, biased for drag reduction).
  • SM batteries:  Battery #7 is in “Cycle” mode; all other batteries (7) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #4 is off line; all other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3928 kg (8660 lb) as of 7/1/04;  [SM(552) + FGB(2772) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 6:51am EDT [= epoch]):

  • Mean altitude — 359.6 km
  • Apogee height — 363.7 km
  • Perigee height — 355.4 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006139
  • Solar Beta Angle — 60.1 deg
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 62 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32291

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.